OpenStudy (anonymous):

How do you find the derivative of x^2 + y^2?

OpenStudy (agreene):

with respect to which variable?

OpenStudy (anonymous):

like for economics we were given U(x,y) = x^2 + y^2 and were asked to find the margin utility with respect to x and y and we have to find partial derivatives but I don't know how to do it... :(

OpenStudy (anonymous):

does that make sense at all??

OpenStudy (agreene):

yes, im trying to remember how to do those partial derivatives in general terms, lol they didnt give you any specific points did they?

OpenStudy (anonymous):

nope, all it says is U(x,y) = x^2 + y^2

OpenStudy (anonymous):

like i have the answer, but i don't know how they got it... they said MUx = 2x and MUy = 2y

OpenStudy (anonymous):

am i just over-thinking the question? lol

OpenStudy (anonymous):

treat y as constant and differentiate then in second x as constant and then differentiate

OpenStudy (agreene):

Oh, okay, thats what I was going to say at first. I thought you were asking for something else. basically, you take the derivative with respect to each variable separately. so d/dx( x^2+y^2) the first goes to 2x (drop the power) the 2nd goes to 0 (no x var so it goes to 0) repeat for y.

OpenStudy (anonymous):

i'm confused :(

OpenStudy (anonymous):

what goes to 0?

OpenStudy (agreene):

the y^2 goes to 0, because there is no x value there. so, by definition of derivatives it will go to 0 when u derive it. Easy mode for derivatives (when you just have polynomials) is to drop the power. so x -> 1 and any constant -> 0

OpenStudy (anonymous):

hmmm... so what is i had something like U(x,y) = (x)(y) + x?

OpenStudy (anonymous):

if i ***

OpenStudy (agreene):

d/dx (x*y+x) = y+1 d/dy (x*y+x) = x

OpenStudy (anonymous):

oh man i dont know why im finding this hard.... dont you have to do the product rule or something? or am i thinking of it too complicating?

OpenStudy (agreene):

No need for product rule, Like Ishaan said, just think of one or the other (x,y) as a constant, or just some number, and do your derivatives... if it's any consolation, lots of people have trouble with this.

OpenStudy (anonymous):

hmmm i think i get it now hahaha... but just to make sure, what if i had something like this then: 3x + xy... would i get this: MUx = 3 + y and MUy = x ????

OpenStudy (agreene):

yes! exactly :)

OpenStudy (anonymous):

woohoo!!!! thanks for your help, i have an assignment due tomorrow and its 2:30 am and i was desperate to get help.... at least now i dont feel completely dumb haha